[Hallc_running] [New Logentry] Hall C RC update: Tuesday, Jan. 18

gaskelld at jlab.org gaskelld at jlab.org
Tue Jan 18 08:25:02 EST 2022

Logentry Text:
Yesterday morning's HMS Q1 issues were quickly resolved, and day shift was able to complete the needed aluminum dummy running and start running on LH2 -  as of 7:15 this morning we have 7298 good e-pi events on LH2. LH2 running at this setting (Q2=3.85, W=3.07, low epsilon, SHMS center) should be completed on day shift today, at which time we'll move on to the SHMS left setting (SHMS theta = 8.39 degrees).

The will be an RC meeting at 10 am this morning:

Zoom link:


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3973446
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220118/6f06f39c/attachment.html>

More information about the Hallc_running mailing list